OpenStudy (anonymous):

Find the hypotenuse of a right triangle whose legs are absolute value of 10 and absolute value of 15 . MEDAL and FAN ASAP help 5 15

OpenStudy (anonymous):

a. absolute value of 5 b. 5 c. 15

OpenStudy (anonymous):

lol i would but i really dont have time for that... sorry

OpenStudy (michele_laino):

hint: ina triangle a side has to be greater than the difference of the other two sides

OpenStudy (lofi):

i think its 5

OpenStudy (anonymous):

ok... so i subract \[\sqrt{15} and \sqrt{10}\]

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (lofi):

yes

OpenStudy (anonymous):

are you sure lol

OpenStudy (anonymous):

no that dont work

OpenStudy (michele_laino):

sorry, \[\sqrt {15} \] is the square root of 15

OpenStudy (anonymous):